### v1

@@ 4 A priori estimates and stability for BSDEs
$\mathbb{E}\!\left[X\ \middle|\ \mathcal{F}_{T-\delta}\right]\ \leq\ C_{2,-}(I)^{1/2}\,\Big(\mathbb{E}^{L}\!\left[X^{2}\,\middle|\,\mathcal{F}_{T-\delta}\right]\Big)^{1/2},$ (4.20)
where the explicit constant $C_{2,-}(I)\ :=\ C_{q}^{*},$ with $q=2$ on interval $I$.
Step 1 (local stability on $I$). Applying Itô’s Lemma to $|\Delta Y|^{2}$ and taking $\mathbb{E}^{L}$ gives
$\mathbb{E}^{L}|\Delta Y_{u}|^{2}+\mathbb{E}^{L}\!\int_{u}^{T}|\Delta Z_{s}|^{2}\,ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\mathbb{E}^{L}\int_{u}^{T}\!\big[\,|\Delta Y_{s}|\,\widetilde{\mathbb{E}}|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\,\big]ds.$ (4.21)
By Corollary 4.5 we have $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$ and the bound depends only on model parameters; hence, for all $u\in I$,
$\mathbb{E}^{L}|\Delta Y_{u}|^{2}+\mathbb{E}^{L}\!\int_{u}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\mathbb{E}^{L}[M_{\delta}],\qquad M_{\delta}:=\int_{T-\delta}^{T}\widetilde{\mathbb{E}}|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\,ds.$ (4.22)
Evaluating (4.22) at $u=T-\delta$,
$\mathbb{E}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,M_{\delta}.$ (4.23)
By Cauchy–Schwarz inequality on the copy space and (4.17), for $s\in I$,
$\widetilde{{\mathbb{E}}}\,|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\;\leq\;\big(\widetilde{{\mathbb{E}}}|\tilde{L}_{s}|^{2}\big)^{1/2}\big(\widetilde{{\mathbb{E}}}|\Delta\tilde{Z}_{s}|^{2}\big)^{1/2}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)^{1/2}\big({\mathbb{E}}|\Delta Z_{s}|^{2}\big)^{1/2}.$
Cauchy–Schwarz in time yields
$M_{\delta}$ $\leq\sqrt{3}\,K_{\tilde{z}}\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)^{1/2}\big({\mathbb{E}}|\Delta Z_{s}|^{2}\big)^{1/2}\,ds$ (4.24)
$\leq\sqrt{3}\,K_{\tilde{z}}\,\Bigg(\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\Bigg)^{\!1/2}\Bigg(\int_{T-\delta}^{T}{\mathbb{E}}|\Delta Z_{s}|^{2}\,ds\Bigg)^{\!1/2}.$
For the first term we simply use BMO energy inequality (kazamaki2006BMO, Sec. 2.1)
$\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\delta+\Big\|\int_{T-\delta}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{T-\delta}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}=:\;C_{\mathrm{energy}},$
where $C_{\mathrm{energy}}<\infty$ and increases with respect to the BMO norm. Hence,
$M_{\delta}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,\Bigg({\mathbb{E}}\!\int_{T-\delta}^{T}\!|\Delta Z_{s}|^{2}\,ds\Bigg)^{\!1/2}.$ (4.25)
Since the right-hand side is deterministic (depends only on $\mathbb{E}$$\cdot$]), the outer expectation ${\mathbb{E}}^{L}$ drops out; thus ${\mathbb{E}}^{L}[M_{\delta}]$ has the same bound. Next, bridge $\mathbb{E}$} and ${\mathbb{E}}^{L}$ by the RHI (4.20) with $X=\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds$:
${\mathbb{E}}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\;\leq\;\sqrt{C_{2,-}(I)}\,\Bigg({\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Bigg)^{\!1/2}.$ (4.26)
Combining (4.25) and (4.26) yields
$M_{\delta}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;\Bigg({\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Bigg)^{\!1/4}.$ (4.27)
To bound the right hand side, pick $u\in I$ and set $N_{u}:=\int_{T-\delta}^{u}\Delta Z_{s}\cdot dW^{L}_{s}$. Then $N]_{T}=\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds$ and the Burkholder–Davis–Gundy (BDG) lower bound at $p=4$ gives $c_{4}^{\downarrow}\,{\mathbb{E}}^{L}\!\big[[N]_{T}^{2}\big]\leq{\mathbb{E}}^{L}|N_{T}|^{4}$. From (4.19),
$N_{T}=X-{\mathbb{E}}^{L}[X|\mathcal{F}_{T-\delta}],\qquad X:=\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\in\mathcal{F}_{T}.$
Thus, Jensen’s inequality and $(|a|+|b|)^{4}\leq 8(|a|^{4}+|b|^{4})$ yield
${\mathbb{E}}^{L}|N_{T}|^{4}\leq 16\,{\mathbb{E}}^{L}|X|^{4}\;\leq\;128\Big({\mathbb{E}}^{L}|\Delta\xi|^{4}+{\mathbb{E}}^{L}\big|\textstyle\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\big|^{4}\Big).$
Hence, using (4.27),
${\mathbb{E}}^{L}\big|\textstyle\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\big|^{4}\;\leq\;M_{\delta}^{4}\;\leq\;9\,K_{\tilde{z}}^{4}\,C_{\mathrm{energy}}^{2}\,C_{2,-}(I)\;{\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}\!|\Delta Z_{s}|^{2}\,ds\Big)^{\!2}\Big].$
Collecting the displays, there exists $C_{\mathrm{BDG}}>0$ (absorbing numeric constants and $c_{4}^{\downarrow}$ only) such that
${\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\;\leq\;C_{\mathrm{BDG}}\Big({\mathbb{E}}^{L}|\Delta\xi|^{4}\;+\;K_{\tilde{z}}^{4}\,C_{\mathrm{energy}}^{2}\,C_{2,-}(I)\;{\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Big).$
By choosing $\delta>0$ small, we can make the local BMO norms of $Z^{1},Z^{2}$ and $\widetilde{{\mathbb{E}}}[L]$ on the slice $I=[T-\delta,T]$ arbitrarily small. Consequently, both the energy constant $C_{\mathrm{energy}}$ (from the BMO energy inequality) and the reverse Hölder constant $C_{2,-}(I)$ are controlled by these norms and $\delta$, hence decrease simultaneously to zero as $\delta\downarrow 0$. In particular, the coefficient in front of ${\mathbb{E}}^{L}\big[(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds)^{2}\big]$ can be made strictly less than $\tfrac{1}{4}$, so we absorb it to the left-hand side and conclude
${\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\;\leq\;4\,C_{\mathrm{BDG}}\;{\mathbb{E}}^{L}|\Delta\xi|^{4}.$
Finally, plug this into (4.27) to get
$M_{\delta}\;\leq\;\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\;\big({\mathbb{E}}^{L}|\Delta\xi|^{4}\big)^{1/4}.$ (4.28)
$\mathbb{H}^{2}$–control of $\Delta Z$ on $I$. Plugging (4.28) into (4.23) gives
$\mathbb{E}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\$$\leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\,(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4}$
$\leq\ \left(2K_{0}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\right)(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4},$
the last inequality using Cauchy–Schwarz and $\Delta\xi\in L^{\infty}(\mathbb{P})$ and thus $\Delta\xi\in L^{\infty}(\mathbb{P}^{L})$ with the same bound $K_{0}$ by Lemma 4.3, and therefore, $\mathbb{E}^{L}|\Delta\xi|^{2}\leq(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/2}\leq 2K_{0}\,(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4}$.
$\mathbb{S}^{2}$–control of $\Delta Y$ on $I$. Recall $X=\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds\in\mathcal{F}_{T}$ from (4.19); for $u\in I$,
$\Delta Y_{u}={\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]-\int_{T-\delta}^{u}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds.$
Hence
$\sup_{u\in I}|\Delta Y_{u}|\ \leq\ \sup_{u\in I}\Big|{\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]\Big|\ +\ \int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\big|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big|\,ds\ =\ \sup_{u\in I}\Big|{\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]\Big|\ +\ M_{\delta}.$
By $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and Doob’s maximal inequality,
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ 8\,{\mathbb{E}}^{L}|X|^{2}\;+2\;{\mathbb{E}}^{L}[M_{\delta}^{2}].$
Using again $(a+b)^{2}\leq 2a^{2}+2b^{2}$,
${\mathbb{E}}^{L}|X|^{2}={\mathbb{E}}^{L}\Big|\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds\Big|^{2}\ \leq\ 2\,{\mathbb{E}}^{L}|\Delta\xi|^{2}\;+2\;{\mathbb{E}}^{L}[M_{\delta}^{2}],$
we obtain
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ 16\,{\mathbb{E}}^{L}|\Delta\xi|^{2}\;+\;18\,{\mathbb{E}}^{L}[M_{\delta}^{2}].$
Invoking the bound (4.28) and $\mathbb{E}^{L}|\Delta\xi|^{2}\leq(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/2}$, we conclude
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ \left(16\;+\;108\,K_{\tilde{z}}^{2}\,C_{\mathrm{energy}}\,\sqrt{C_{2,-}(I)}\,\sqrt{C_{\mathrm{BDG}}}\right)\big({\mathbb{E}}^{L}|\Delta\xi|^{4}\big)^{1/2}.$ (4.29)
We arrive at the local estimate $({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/2}\leq 2K_{0}({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/4}$:
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\;+\;{\mathbb{E}}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ C_{\mathrm{loc}}(I)\,({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/4},$ (4.30)
with the constant $C_{\mathrm{loc}}$ depending on $\delta$, $K_{0}$, $K_{z}$, $K_{\tilde{z}}$ and on the BMO/RHI/BDG constants.
Step 2 (global stability). Let $0=t_{0}<t_{1}<\cdots<t_{N}=T$ with $t_{k+1}-t_{k}=\delta$ (small enough so the local estimates hold on each slice $I_{k}:=[t_{k},t_{k+1}]$). By Proposition 4.4 we have $\Delta Y_{t_{k+1}}\in L^{\infty}(\mathbb{P})$. Since essential suprema are invariant under equivalent changes of measure (Lemma 4.3), in particular, $\Delta\xi$ and $\Delta Y_{t_{k+1}}$ belong to $L^{\infty}$ under both $\mathbb{P}$ and $\mathbb{P}^{L}$. Hence (4.30) applies on every $I_{k}$ with terminal datum $\Delta Y_{t_{k+1}}\in L^{\infty}(\mathbb{P}^{L})$:
$\mathbb{E}^{L}\!\Big[\sup_{u\in I_{k}}|\Delta Y_{u}|^{2}\Big]\;+\;\mathbb{E}^{L}\!\int_{I_{k}}|\Delta Z_{s}|^{2}\,ds\ \leq\ C_{\mathrm{loc}}(I_{k})\;\Big(\mathbb{E}^{L}|\Delta Y_{t_{k+1}}|^{4}\Big)^{1/4}.$ (4.31)
Repeating Step 1 on $I_{k}$ with terminal $\Delta Y_{t_{k+1}}$ under $\mathbb{P}^{L}$ read
$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW_{s}^{L},\qquad u\in I_{k}.$
Evaluating at $u=t_{k}$ and setting
$X_{k}:=\Delta Y_{t_{k+1}}+\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds,\qquad M_{\delta,k}:=\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big|\,ds,$
we get
$\Delta Y_{t_{k}}={\mathbb{E}}^{L}\!\big[X_{k}\,\big|\,\mathcal{F}_{t_{k}}\big]-\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds.$
Using $(a+b)^{4}\leq 8(a^{4}+b^{4})$ and Jensen’s inequality,
${\mathbb{E}}^{L}\!|\Delta Y_{t_{k}}|^{4}$ $\leq 8\,{\mathbb{E}}^{L}|X_{k}|^{4}+8\,{\mathbb{E}}^{L}\Big|\textstyle\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\Big|^{4}\leq 8\,{\mathbb{E}}^{L}|X_{k}|^{4}+8\,M_{\delta,k}^{4}$
$\leq 8(8\,{\mathbb{E}}^{L}|\Delta Y_{t_{k+1}}|^{4}+8\,M_{\delta,k}^{4})+8\,M_{\delta,k}^{4}$
$=64\,{\mathbb{E}}^{L}\!|\Delta Y_{t_{k+1}}|^{4}\;+\;72\,M_{\delta,k}^{4}$
By repeating the bound for $M_{\delta}$ derived on a single slice (with $\Delta\xi$ replaced by the terminal datum $\Delta Y_{t_{k+1}}$), there exists a constant $C_{\delta}(I_{k})$ (depending only on $K_{z},K_{\tilde{z}},K_{0}$ and the BMO energy/RHI/BDG constants) such that
$M_{\delta,k}\;\leq\;C_{\delta}(I_{k})\,\Big({\mathbb{E}}^{L}\big[|\Delta Y_{t_{k+1}}|^{4}\big]\Big)^{1/4}.$
Therefore
$\Big({\mathbb{E}}^{L}|\Delta Y_{t_{k}}|^{4}\Big)^{1/4}\;\leq\;\alpha_{\delta}(I_{k})\;\Big({\mathbb{E}}^{L}\!\big[|\Delta Y_{t_{k+1}}|^{4}\big]\Big)^{1/4},\qquad\alpha_{\delta}(I_{k}):=\big(64+72\,C_{\delta}(I_{k})^{4}\big)^{1/4}.$ (4.32)
Iterating gives
$\Big(\mathbb{E}^{L}|\Delta Y_{t_{k+1}}|^{4}\Big)^{1/4}\ \leq\ \prod\limits_{i=k}^{N-k-1}\alpha_{\delta}(I_{i+1})\;\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}.$ (4.33)
Summing (4.31) over $k$ and using (4.33),
$\mathbb{E}^{L}\!\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big]+\mathbb{E}^{L}\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$ $\leq\sum_{k=0}^{N-1}\left(\mathbb{E}^{L}\!\Big[\sup_{u\in I_{k}}|\Delta Y_{u}|^{2}\Big]+\mathbb{E}^{L}\!\int_{I_{k}}|\Delta Z_{s}|^{2}\,ds\right)$
$\leq\left(\sum_{k=0}^{N-1}C_{\mathrm{loc}}(I_{k})\,\prod\limits_{i=k}^{N-k-1}\alpha_{\delta}(I_{i+1})\right)\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}$
$=:C_{\mathrm{global}}\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}.$
Finally, the Girsanov kernel is a BMO martingale, the RHI transfers this global $\mathbb{P}^{L}$ bound back to $\mathbb{P}$, up to a multiplicative factor. This completes the proof. ∎
Remark 4.7.
Theorem 4.6 provides a stability estimate in $\mathbb{S}^{2}\times\mathbb{H}^{2}$. The proof is based on a global linearisation argument and can be generalized to obtain analogous $L^{p}$ stability estimates (for $p>2$) by similar argument. We do not pursue this level of generality here, as the $\mathbb{S}^{2}\times\mathbb{H}^{2}$ control is already sufficient for our purposes in the sense that the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is continuously embedded in $\mathbb{S}^{2}\times\mathbb{H}^{2}$.
In particular, in the existence proof for the main result below, we construct approximating solutions and show that the sequence is Cauchy in $\mathbb{S}^{2}\times\mathbb{H}^{2}$ by applying Theorem 4.6. The limit in $\mathbb{S}^{2}\times\mathbb{H}^{2}$ then automatically belongs to $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
5 Proof: existence and uniqueness of solutions

### v2

@@ 4 A priori estimates and stability for BSDEs
$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]+\varepsilon\sqrt{T}\,\left(\int_{u}^{T}\mathbb{E}|\Delta Z_{s}|^{2}\,ds\right)^{1/2},\qquad u\in[0,T].$
By BMO energy inequality (kazamaki2006BMO, Sec. 2.1) and Corollary 4.5, the last term is bounded by a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ since:
$\int_{u}^{T}\Big({\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\Big\|\int_{0}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{0}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}.$
Therefore, letting $\varepsilon\to 0$ yields
$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]=(\Lambda_{u})^{-1}\,\mathbb{E}[\Lambda_{T}|\Delta\xi|\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$ (4.22)
Now fix $q\in(1,q^{*})$ and let $p$ be its conjugate exponent. By the reverse Hölder inequality and conditional Hölder’s inequality,
$|\Delta Y_{u}|^{p}$ $\leq(\Lambda_{u})^{-p}\,\mathbb{E}[(\Lambda_{T})^{q}\,|\,\mathcal{F}_{u}]^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}]$
$\leq(C_{q}^{*})^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$
By Doob’s maximal inequality,
$\mathbb{E}\Big(\sup_{u\in[0,T]}|\Delta Y_{u}|^{p}\Big)\ \leq\ C_{Y}(p)\,\mathbb{E}|\Delta\xi|^{p},\qquad p\geq p^{*},$ (4.23)
where $p^{*}$ is the conjugate exponent of $q^{*}$ and $C_{Y}(p):=\Big(\frac{p}{p-1}\Big)^{p}(C_{q}^{*})^{p/q}$. They both depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ as we discussed in the Preliminary step.
$\mathbb{H}^{p}$–control of $\Delta Z$. Fix $p>p^{*}$, and applying Itô’s formula to $|\Delta Y|^{2}$ on $[0,T]$. we obtain
$|\Delta Y_{0}|^{2}+\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds=|\Delta\xi|^{2}+2\int_{0}^{T}\Delta Y_{s}\,\widetilde{\mathbb{E}}[\Delta g(s,Z_{s},\tilde{Z}_{s})]\,ds-2\int_{0}^{T}\Delta Y_{s}\,Z_{s}\cdot dW_{s}$
Using the same Fefferman/Cauchy–Schwarz argument as above and $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$, we obtain
$2\int_{0}^{T}\Delta Y_{s}\,\widetilde{{\mathbb{E}}}\!\big[\Delta g(s,Z_{s},\tilde{Z}_{s})\big]\,ds$ $\leq 2K_{z}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)\,|\Delta Y_{s}|\,|\Delta Z_{s}|\,ds$
$\quad+2K_{\tilde{z}}\int_{0}^{T}|\Delta Y_{s}|\,\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds$
$\leq\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big(2K_{z}^{2}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds\Big)+\frac{1}{2}\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$
$\quad+2K_{\tilde{z}}\|\Delta Y\|_{\mathbb{S}^{\infty}}\int_{0}^{T}\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds.$
The last term vanishes in the limit $\varepsilon\to 0$ by the same slicing argument as in (4.21). Next, let $K:=\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds$. Then
${\mathbb{E}}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\,K^{p/2}\Big]\leq\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[K^{p}\big]\Big)^{1/2}.$ (4.24)
If $\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}$ is a BMO martingale ($i=1,2$), the BMO energy inequality (see (kazamaki2006BMO, Sec. 2.1)) gives
${\mathbb{E}}\Big[\Big(\int_{0}^{T}|Z_{s}^{i}|^{2}ds\Big)^{p}\Big]\leq p!\,\Big\|\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}.$
Therefore,
${\mathbb{E}}[K^{p}]\leq 3^{p-1}\Big((3T)^{p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}\Big)=:C_{\mathrm{energy}}(p).$ (4.25)
Lastly, with the help of BDG inequality and Cauchy–Schwarz, we have
${\mathbb{E}}\Big|\int_{0}^{T}(\Delta Y_{s})\Delta Z_{s}\cdot dW_{s}\Big|^{p/2}$ $\leq C_{\mathrm{BDG}}\,{\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Y_{s}|^{2}|\Delta Z_{s}|^{2}\,ds\Big)^{p/4}\Big]$
$\leq C_{\mathrm{BDG}}\,\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[(\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds)^{p/2}\big]\Big)^{1/2}.$ (4.26)
Putting (4.24), (4.25) and (4) together, we obtain
${\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}{\mathbb{E}}[|\Delta\xi|^{p}]+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\Big({\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}]\Big)^{1/2}$
$\quad+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,{\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}],$
From previous step (4.23), there exist constants $C_{Y}(p),C_{Y}(2p)$ such that
$\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\Big]\leq C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big],\qquad\Big(\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\Big]\Big)^{1/2}\leq\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
Substituting these bounds into the previous inequality, we obtain for all $p\geq p^{*}$,
$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}\mathbb{E}\big[|\Delta\xi|^{p}\big]+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big]$
$\quad+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
Finally, since $\mathbb{E}|\Delta\xi|^{p}\leq\big(\mathbb{E}|\Delta\xi|^{2p}\big)^{1/2}$, we may absorb the first two terms into the last one and conclude that
$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq C_{Z}(p)\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2},$
where
$C_{Z}(p):=2^{\frac{3p}{2}-1}+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}.$
Combing the stability estimate (4.23) completes the proof. ∎
5 Proof: existence and uniqueness